OpenStudy (anonymous):

Find the sum of the infinite geometric series in which a1=8 and r= -1/2

OpenStudy (anonymous):

use the formula sum = a ----------- 1-r

OpenStudy (anonymous):

is it k? or should i give u the ans

OpenStudy (anonymous):

Give me the answer.

OpenStudy (anonymous):

sry not 16..........it should be 16/3

OpenStudy (anonymous):

put limit n tends infi and u'l see myne formula has nothin wrong......no offence

OpenStudy (anonymous):

1/2 to the power n where n is infi will only make it smaller

OpenStudy (anonymous):

jose ur goin in wrong direction

OpenStudy (anonymous):

cmon answer me (1/2)^infinity is smaller or larger

OpenStudy (anonymous):

then 1/2

OpenStudy (anonymous):

thats not abt getting it determined but u can say that 1/2^inf must be very small..........so it can be done if u neglect that value

OpenStudy (anonymous):

but its very important to note my formula works only when r<1........

OpenStudy (anonymous):

the same way inf-1 is inf itself ..........:) don't get frustrated on me....... try to understand

OpenStudy (anonymous):

=8 / (3/2) = 8 x2/3 = 16/3 thats it, end of story

OpenStudy (anonymous):

thanks elecengineer

OpenStudy (anonymous):

the sum of a GP is \[\frac{a(1-r^n)}{1-r}\]

OpenStudy (anonymous):

if the absolute value of r is less then 1 , then (1-r^n) ---> 1 , as n-> infinity

OpenStudy (anonymous):

so you are multiplying by a factor of 1, and thus why it disappears

OpenStudy (anonymous):

r^n gets very small if |r|<1 think of 0.01^1000 , thats 0.01 x 0.01 x 0.01 ... etc 1000 times, it quickly approaches zero therefore 1-r^n appraoches 1

OpenStudy (anonymous):

elecengineer u r very good at explaining :)

OpenStudy (anonymous):

likewise if you had (-0.01)^999 ( an odd power so its not even ) then you would get -0.01 x -0.01x ... etc

OpenStudy (anonymous):

that would give a very small negative answer, so 1-r^n would be just a tiny bit bigger than 1 , but it would still approach 1

OpenStudy (anonymous):

All the above is also explained here: http://en.wikipedia.org/wiki/Geometric_series Scroll to "formula" section.

OpenStudy (anonymous):

or be very close to 1, so in the limit it is neglected

OpenStudy (anonymous):

I stated a wrong equation which lead me to a wrong answer. Sorry for the inconvenience. LOL. I really should start proofreading my works. BTW, thanks for the corrections. Sorry Ishaan. :))